A market for baby bottles has the following supply and demand functions

qS = −6 + 3p

qD = 14 − 2p

a. Now, suppose a per unit tax of 5 were charged to the buyer. What are the equilibrium quantity, price paid by the buyer, and price received by the seller?

b. How much tax revenue is raised? How much of that tax burden is borne by the buyer?

c. Calculate the Consumer Surplus, Producer Surplus, Total Welfare Level, and Deadweight Loss associated with the tax.
Answer #1

s Qs = -6 +3P Qd = 14-2P In Equilibrium Qd = QS lu-ap =-6+3P 14+6 =3ptap 20 =SP P = 20 P =64 substituting pinto as equation Du-2P =- 6+3 p 4+6 = 3PtQP lo - SP P = 10 seller Receive P = $2. Subst stituting pinto as equation Qs =-6+3P =-6 + 3x2 =-6 +6

Buyra on buyrer = – 4 = $ 3 Tax burden on seller = u-2 Price u 6 8 10 12 14 16 18 202, (0) Consumer surplus decrease by Areac.s =$9 Decrease in productos surplus by Area (B) = area of A = 1 XB XH =o.sx 6x (4-2) = 3x2 DWL = Area of A (A+B) x 8 x H -0